Dear all,
please find below a short summary of the activities related to the AO
system.


People at LBTO:


R. Briguglio



Summary
Operations began at 1:40UT, with the shell successfully set. The gap was inspected before setting the shell and it was found on its nominal positions. However, at later inspections the 60um ear was detected for several minutes. The system worked until 12:27UT when the shell was ripped at the end of operations. A schematic log of working time may be summarized as follows:
1:40UT shell set
04:58 stop, current alarm for act#530
05:00 shell set again
07:03 stop: long series of errors (see log for details)
08:25 shell set successfully again
12:27 end of operations.
Act#45 was removed from current control and this allowed to keep on working. The actuators that gave at least an error are reported below:
Currents errors: #530, #45, #4, #537. Position error: #4. Act #240, #483 and#4 gave currents warning, while #19 suffered for a position RMS warning.
At approx 7:40UT, while trying to recover the shell from a series or errors, the telescope was moved by 180° in azimut, pointing in a direction opposite to the wind, so that the measured wind speed slew down from 6m/s to 1.5m/s. At that point the shell was set without problems.

The cooling system suffered by an error detected at approx 08UT (a valve that was found stucked). The cooling fluid temperature increased from 2°C to 9°C: the failure was fixed by resetting remotely the valve.

A more detailed log may be found here:
TelLogLBT672a20100405
also some images are attached, including a .doc file where temperature, windspeed, flux rate are plotted versus time, for the whole working time.


Ciao
Runa

-- ArmandoRiccardi - 06 Apr 2010
Topic revision: r1 - 06 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback